the sum of the digits of a two digit number is 7. if the digits are reversed, the newly formed number is 9 larger than the original number. Find the original number.
Answer by jorel1380(3719)

x+y=7
10x+y=10y+x-9
-------------
y=7-x
10x+(7-x)=10(7-x)+x-9
9x+7=70-10x+x-9
""=61-9x
18x=54
x=3
y=4
The original number is 34.